A low-noise transistor for use in computing products is being developed. It is claimed that the mean noise level will be below the 2.5-dB level of products currently in use. It is believed that the noise level is approximately normal with a standard deviation of 0.8.

Find a 95% CI for the mean noise level of the population.
Answer #1

Answer:

Given Data

The 95% Confidence interval

=\bar{X}\pm z_{\alpha/2}\frac{\sigma}{\sqrt{n}}\\1.8 \pm 1.96*\frac{0.8}{\sqrt{16}}

1.8 \pm 1.96(0.2)

2.192 < \mu < 1.408

****Please like it...

It is important to me..

Thank you for supporting me....